At a car dealership,3/10 of the vehicles are trucks, 1/5 are sports cars, and the rest are vans. If there are 90 trucks, how many vans are there?
Answer by Mathtut(3670)

the number of vans is 1-3/10-2/10=5/10 of all vehicles
:
x=total number of vehicles
:
3/10x=90 --->3x=900 --->x=300 trucks
:
vans =5/10(300)=150 # of vans
:
: